# Data-Locality-Aware Task Assignment and Scheduling for Distributed Job Executions

### v3

@@ III-B Water-Filling
Solving a group of integer programs might still be time-consuming for large-scale distributed computing systems. To improve efficiency, we propose another task assignment algorithm called water-filling (WF), extended from [7], which assigns tasks one group at a time in an approximate manner.
@@ III-B1 Algorithm Design
We denote by $b_{m}^{c}(k)$ the estimated busy time of server $m$ after assigning the tasks in group $k$. By definition, for each $m\in\mathcal{M}$, we have
$b_{m}^{c}(0):=b_{m}^{c}\leq b_{m}^{c}(1)\leq\cdots\leq b_{m}^{c}\left(K_{c}\right).$ (10)
For each $k\in\mathcal{K}_{c}$, WF allocates $\max\{\xi_{k}-b_{m}^{c}(k-1),0\}$ time slots on each available server $m\in\mathcal{S}_{c}^{k}$, where $\xi_{k}$ is the minimal integer such that
$\sum_{m\in\mathcal{S}_{c}^{k}}\left(\max\left\{\xi_{k}-b_{m}^{c}(k-1),0\right\}\right)\cdot\mu_{m}^{c}\geq|\mathcal{T}_{c}^{k}|.$ (11)
The calculation of $\xi_{k}$ is similar to $x_{k}$ in (9). It is the estimated time for all the tasks in group $k$ to be processed by server $m$, given the assignment of tasks in $\mathcal{T}_{c}^{1},\ldots,\mathcal{T}_{c}^{k-1}$. We call $m\in\mathcal{S}_{c}^{k}$ a participating server if $\xi_{k}-b_{m}^{c}(k-1)>0$. We then allocate $(\xi_{k}-b_{m}^{c}(k-1))\cdot\mu_{m}^{c}$ tasks in group $k$ to each participating server $m$ (or all the remaining tasks in group $k$ if server $m$ is the last participating server). After that, we update $b_{m}^{c}(k)$ by
$b_{m}^{c}(k)=\max\left\{b_{m}^{c}(k-1),\xi_{k}\right\}$ (12)
for each $m\in\mathcal{S}_{c}^{k}$. WF repeats this procedure for each task group sequentially, performing like filling water, ensuring that the participating servers have the same estimated busy time. Fig. 2 illustrates the idea of water-filling.
@@
The pseudocode of WF is presented in Algorithm 2. The complexity of WF is $\mathcal{O}\left(\sum_{k\in\mathcal{K}_{c}}|\mathcal{S}_{c}^{k}|\log(|\mathcal{T}_{c}^{k}|)\right)$, since for each group $k$, $\xi_{k}$ can be found through a binary search with an overhead of $\mathcal{O}(\log(|\mathcal{T}_{c}^{k}|))$, and WF needs to walk through all the available servers to verify that the searched $\xi_{k}$ satisfies (11). Note that the complexity of WF can also be written as $\mathcal{O}\left(K_{c}\cdot M\cdot\log|\mathcal{T}_{c}|\right)$.
@@
2       for each server $m\in\bigcup_{k\in\mathcal{K}_{c}}\mathcal{S}_{c}^{k}$ do
@@ 4             $b_{m}^{c}(0)\leftarrow b_{m}^{c}$
5      for each task group $k\in\mathcal{K}_{c}$ do
@@
11                         Assign $\left(\xi_{k}-b_{m}^{c}(k-1)\right)\cdot\mu_{m}^{c}$ group-$k$ tasks to server $m$
@@ III-B2 Approximation Analysis
In this section, we present the performance analysis of WF. The approximation factor of WF against the optimum is defined as
@@
where $I:=I\big{(}c,\{b_{m}^{c}\}_{m}\big{)}$ is an arrival instance of a new job $c$ and the initial estimated busy time of each server just before job $c$’s arrival. $\text{OPT}(I)$ and $\text{WF}(I)$ are respectively the maximum estimated busy times of participating servers after the tasks of job $c$ are assigned by the optimal algorithm and WF.
Figure 3: A visualization of the assignment made by WF for the constructed instance.
@@
The WF-to-optimal ratio is 1 whenever $\mathcal{T}_{c}$ has only one task group or $\mathcal{S}_{c}^{k}\cap\mathcal{S}_{c}^{k^{\prime}}=\emptyset$ holds for any $k\neq k^{\prime}$. Thus, to get a worst-case ratio, we construct an instance $I$ as illustrated in Fig. 3, where the available servers of different task groups overlap.
For simplicity, we assume $\mu_{m}^{c}=1$ and $b_{m}^{c}=0$ for each server $m$. For each group $k$, the number of its available servers is given by $|\mathcal{S}_{c}^{k}|=\sum_{k^{\prime}=1}^{K_{c}-k+1}\theta^{k^{\prime}}$, where $\theta\geq 2$ is an integer. Moreover,
@@
i.e., the available servers of a higher-indexed group are a subset of those of a lower-indexed group. The number of tasks in group $k$ is given by $|\mathcal{T}_{c}^{k}|=\theta\cdot|\mathcal{S}_{c}^{k}|$. The task assignment made by WF is shown in Fig. 3, where $\theta$ time slots are required to process the tasks in each group.
Figure 4: A visualization of the assignment made by OPT for the constructed instance.
By contrast, in OPT, for each $k<K_{c}$, the tasks in group $k$ are assigned to a subset of its available servers $\mathcal{S}_{c}^{k}\backslash\mathcal{S}_{c}^{k+1}$ only, as illustrated in Fig. 4. Then, the number of time slots needed to process all the tasks in group $k$ is
$\Bigg{\lceil}\frac{\theta\cdot\sum_{k^{\prime}=1}^{K_{c}-k+1}\theta^{k^{\prime}}}{\theta^{K_{c}-k+1}}\Bigg{\rceil}=\theta+2.$ (15)
Additionally, the number of time slots needed to process all the tasks in group $K_{c}$ is $\theta$. Hence, by OPT, the completion time of job $c$ is $\textsc{OPT}(I)=\theta+2$.
Thus, the WF-to-optimal ratio is
$\frac{\textsc{WF}(I)}{\textsc{OPT}(I)}=\frac{K_{c}\cdot\theta}{\theta+2},$ (16)
which can be made arbitrarily close to $K_{c}$ as $\theta$ approaches infinity. ∎
Theorem 1 shows that $K_{c}$ is a lower bound of $\alpha_{\textsc{WF}}$. In the following, we prove that $K_{c}$ is also an upper bound of $\alpha_{\textsc{WF}}$.

### v4

@@ III-B Water-Filling
Solving a group of integer programs can still be time-consuming for large-scale services environments. To improve efficiency, we propose an alternative task assignment algorithm called Water-Filling (WF), which is extended from [9]. WF assigns tasks to jobs sequentially, one group at a time, using an approximate approach to balance computational efficiency with performance.
@@ III-B1 Algorithm Design
WF dynamically estimates server workloads as tasks are assigned. Let $b_{m}^{c}(k)$ denote the estimated busy time of server $m$ after assigning tasks from group $k$. Initially, for each server $m\in\mathcal{M}$, the busy time is:
$b_{m}^{c}(0):=b_{m}^{c},$ (10)
and it increases monotonically as task groups are assigned:
$b_{m}^{c}(0)\leq b_{m}^{c}(1)\leq\cdots\leq b_{m}^{c}\left(K_{c}\right).$ (11)
For each task group $k$, WF determines the minimum time $\xi_{k}$ required to process all tasks in the group across its available servers. Specifically, $\xi_{k}$ is the smallest integer satisfying:
$\sum_{m\in\mathcal{S}_{c}^{k}}\left(\max\left\{\xi_{k}-b_{m}^{c}(k-1),0\right\}\right)\cdot\mu_{m}^{c}\geq|\mathcal{T}_{c}^{k}|,$ (12)
where $\mathcal{S}_{c}^{k}$ is the set of servers available for group $k$, and $b_{m}^{c}(k-1)$ represents the server’s busy time after the previous task group assignment. The calculation of $\xi_{k}$ is analogous to $x_{k}$ in (9), representing the estimated time needed to process all tasks in the group. A server $m\in\mathcal{S}_{c}^{k}$ is considered a participating server if $\xi_{k}-b_{m}^{c}(k-1)>0$. We then allocate $(\xi_{k}-b_{m}^{c}(k-1))\cdot\mu_{m}^{c}$ tasks in group $k$ to each participating server $m$ (or all the remaining tasks in group $k$ if server $m$ is the last participating server). After the assignment, the busy time of each server is updated as:
$b_{m}^{c}(k)=\max\left\{b_{m}^{c}(k-1),\xi_{k}\right\}.$ (13)
This process, repeated for each task group, simulates a water-filling mechanism where the participating servers’ workloads are gradually equalized. Fig. 2 illustrates the water-filling concept for task assignment.
@@
The pseudocode of WF is presented in Algorithm 2. Its computational complexity is $\mathcal{O}\left(\sum_{k\in\mathcal{K}_{c}}|\mathcal{S}_{c}^{k}|\log(|\mathcal{S}_{c}^{k}|)\right)$, as $\xi_{k}$ can be calculated via binary search with $\mathcal{O}(\log(|\mathcal{S}_{c}^{k}|))$ iterations. This complexity can also be expressed as $\mathcal{O}(K_{c}\cdot M\cdot\log M)$.
Input: Online arriving jobs $1,2,...,c,...$, server capacities $\{\mu_{m}^{c}\}_{m,c}$
Output: Assignment solution for each task of each arriving job
@@
2 for each $m\in\bigcup_{k\in\mathcal{K}_{c}}\mathcal{S}_{c}^{k}$ do
@@ 4    $b_{m}^{c}(0)\leftarrow b_{m}^{c}$
5 for each $k\in\mathcal{K}_{c}$ do
@@
[…]
@@ III-B2 Approximation Analysis
In this section, we analyze the performance of WF. The approximation factor of WF, compared to the optimal algorithm, is defined as:
@@
where $I:=I\big{(}c,\{b_{m}^{c}\}_{m}\big{)}$ represents the arrival instance of a new job $c$ and the initial estimated busy times of all servers prior to job $c$’s arrival. Here, $\text{OPT}(I)$ and $\text{WF}(I)$ denote the maximum estimated busy times of participating servers after assigning job $c$’s tasks using the optimal algorithm and WF, respectively.
Figure 3: Visualization of the task assignment by WF for a constructed instance.
@@
The WF-to-optimal ratio is 1 under favorable conditions, such as when $\mathcal{T}_{c}$ consists of only one task group or when $\mathcal{S}_{c}^{k}\cap\mathcal{S}_{c}^{k’}=\emptyset$ for all $k\neq k’$ ’. However, to demonstrate the worst-case ratio, we construct a specific instance $I$, illustrated in Fig. 3, where the available servers of different task groups overlap.
For simplicity, let $\mu_{m}^{c}=1$ and $b_{m}^{c}=0$ for all servers $m$. For each task group $k$, the number of its available servers is defined as:
$|\mathcal{S}_{c}^{k}|=\sum_{k^{\prime}=1}^{K_{c}-k+1}x^{k^{\prime}},$ (15)
where $x\geq 2$ is an integer. Furthermore, the available servers follow a hierarchical structure:
@@
meaning the servers available for a higher-indexed group are a subset of those available for a lower-indexed group. Each task group $k$ contains
$|\mathcal{T}_{c}^{k}|=x\cdot|\mathcal{S}_{c}^{k}|$
tasks. In WF, tasks in each group are assigned sequentially, and as shown in Fig. 3, $x$ time slots are required to process all tasks in each group.
Figure 4: Visualization of the task assignment by OPT for the constructed instance.
In the optimal assignment (OPT), illustrated in Fig. 4, tasks from group $k$ ($k<K_{c}$) are assigned to a subset of its servers $\mathcal{S}_{c}^{k}\setminus\mathcal{S}_{c}^{k+1}$. The number of time slots required to process all tasks in group $k$ is:
$\Bigg{\lceil}\frac{x\cdot\sum_{i=1}^{K_{c}-k+1}x^{i}}{x^{K_{c}-k+1}}\Bigg{\rceil}=x+2.$ (17)
For group $K_{c}$, the number of time slots required remains $x$. Therefore, the total completion time for job $c$ using OPT is:
$\textsc{OPT}(I)=\max\{x+2,x\}=x+2.$ (18)
The approximation factor for this instance is:
$\frac{\textsc{WF}(I)}{\textsc{OPT}(I)}=\frac{K_{c}\cdot x}{x+2},$ (19)
As $x\to\infty$, the ratio approaches $K_{c}$, demonstrating that the worst-case approximation factor of WF is at least $K_{c}$. ∎
Theorem 1 establishes that $K_{c}$, the number of task groups in job $c$, is a lower bound for the approximation factor $\alpha_{\textsc{WF}}$ of WF. We now demonstrate that $K_{c}$ is also an upper bound for $\alpha_{\textsc{WF}}$, thereby tightly characterizing the approximation factor of WF.